OpenStudy (anonymous):

how does the integral of dx/(xsqrt4x^2-1) become du/(usqrtu^2-1) when u=2x? specifically the 4x^2 becoming u^2

zepdrix (zepdrix):

Bring the 4 into the square with the x, \[\Large\rm \int\limits\limits \frac{dx}{x \sqrt{4x^2-1}}=\int\limits\limits \frac{dx}{x \sqrt{(2x)^2-1}}\]

zepdrix (zepdrix):

Then we need to use these three pieces for our substitution: \[\Large\rm u=2x, \qquad\qquad x=\frac{1}{2}u\qquad\qquad du=2dx\]

zepdrix (zepdrix):

Would I should have written that third piece as: \(\Large\rm \dfrac{1}{2}du=dx\) Plugging them in gives us: \[\Large\rm \int\limits \frac{\frac{1}{2}du}{\frac{1}{2}u \sqrt{(u)^2-1}}\]

zepdrix (zepdrix):

Woops* not would.. blah

OpenStudy (anonymous):

thank you again!